If you were looking at the healthy fitness rate on patients and your current rate was 92% and this was out of the last 1550 patients that you deal with. If you wanted your rate to increase to 96% for the next 1550, how many more would you have to put into the list?  What would be the step by step process?
Answer #1

Current rate = 92% = 0.92

Number of healthy patients = 0.92 * 1550 = 1426

If you wanted your rate to increase to 96% for the next 1550, the target for number of healthy patients = 0.96 * 1550 = 1488

Number of more patients to put into the list = 1488 - 1426 = 62